# Pairs of Clark Unitary Operators on the Bidisk and their Taylor Joint Spectra

### v1

@@ 2. A Simple Example
The family $\{\sigma_{\alpha}\}_{\alpha\in\mathbb{T}}$ of Clark measures associated with $\phi$ can readily be identified, see eg. [20, Example 4.2]. Namely, we have the pairing
$\int_{\mathbb{T}^{2}}fd\sigma_{\alpha}=\int_{\mathbb{T}}f(\zeta_{1},\alpha\bar{\zeta}_{1})|d\zeta_{1}|,\quad f\in C(\mathbb{T}^{2}).$
In other words, $\mathrm{supp}\,\sigma_{\alpha}=\{(\zeta_{1},\alpha\bar{\zeta}_{1})\colon\zeta_{1}\in\mathbb{T}\}$. As is explained in [20] (or more generally as guaranteed by [4, Theorem 3.3]), there is a surjective isometry $J_{\alpha}\colon K_{\phi}\to L^{2}(\sigma_{\alpha})$ for each value of $\alpha$. (By [9, Theorem 1.3], this isometry acts by evaluation of $f\in K_{\phi}$ on $\mathrm{supp}\,\sigma_{\alpha}$, which is well-defined for $\sigma_{\alpha}$-a.e. $\zeta\in\mathbb{T}^{2}$, but here this can be checked directly.) Hence, on $\mathrm{supp}\,\sigma_{\alpha}$,
$(J_{\alpha}U^{1}_{\alpha})(z_{1}^{k})=\zeta_{1}\cdot\zeta_{1}^{k}\quad\textrm{and}\quad(J_{\alpha}U^{1}_{\alpha})(z_{2}^{l})=\zeta_{1}\cdot\alpha\bar{\zeta}_{1}^{l},$
so that $J_{\alpha}U^{1}_{\alpha}=M_{\zeta_{1}}$, where $M_{\zeta_{1}}\colon L^{2}(\mu_{\alpha})\to L^{2}(\mu_{\alpha})$ is the unitary furnished by multiplication by the coordinate function $\zeta_{1}$. Similarly,
$(J_{\alpha}U^{2}_{\alpha})(z_{1}^{k})=\alpha\bar{\zeta}_{1}\cdot\zeta_{1}^{k}\quad\textrm{and}\quad(J_{\alpha}U^{2}_{\alpha})(z_{2}^{l})=\alpha\bar{\zeta}_{1}\cdot\alpha\bar{\zeta}_{1}^{l},$
which implies that $J_{\alpha}U^{2}_{\alpha}=M_{\alpha\bar{\zeta}_{1}}=M_{\zeta_{2}}$.
Finally, one can check directly (or refer to Theorem 5.5 below) that the Taylor joint spectrum of $(U^{1}_{\alpha},U^{2}_{\alpha})$ (see Subsection 3.3 for a definition) is
@@ 3.1. Clark Theory in One Variable
Recall that each inner function $\phi\colon\mathbb{D}\to\mathbb{C}$ gives rise to a family of Clark measures $\{\sigma_{\alpha}\}_{\alpha\in\mathbb{T}}$ via (1.1); each $\sigma_{\alpha}$ is singular with respect to Lebesgue measure on the unit circle, and moreover $\sigma_{\alpha_{1}}\perp\sigma_{\alpha_{2}}$ for distinct $\alpha_{j}$. Clark measures can be identified as spectral measures associated with certain unitary operators acting on the model space $K_{\phi}=H^{2}(\mathbb{D})\ominus\phi H^{2}(\mathbb{D})$. Namely, while the compressed shift $S_{\phi}\colon K_{\phi}\to K_{\phi}$ is not generally unitary (see eg. [22, Chapter 9]), they do admit one-dimensional perturbations to unitaries. More specifically, for each $\alpha\in\mathbb{T}$, we set
$U_{\alpha}f:=S_{\phi}f+\frac{\alpha}{1-\overline{\phi(0)}\alpha}\langle f,Ck_{0}\rangle k_{0},\quad f\in K_{\phi},$ (3.1)
where $k_{0}$ is the reproducing kernel for $K_{\phi}$ at the origin, and $Ck_{0}$ is its conjugation. If $\phi(0)=0$, then $k_{0}=1$ and (3.1) simplifies to
@@
and then extending to $K_{\phi}$ by linearity. Then we have the following theorem.
@@
See [22, Chapter 8, Chapter 11] for a detailed exposition, including proofs of the above statements.
@@
Then the associated model space $K_{\phi}$ has $\dim K_{\phi}=\deg\phi$ (see [22, Proposition 5.16]), hence is finite-dimensional, and is spanned by reproducing kernels and their derivatives. (In fact, $K_{\phi}$ is finite dimensional precisely when $\phi$ is a finite Blaschke product.)
@@
$\sigma_{\alpha}=\sum_{\mathcal{C}_{\alpha}}\frac{1}{|\phi^{\prime}(\zeta)|}\delta_{\zeta},$ (3.2)
where $\mathcal{C}_{\alpha}=\{\zeta\in\mathbb{T}\colon\phi(\zeta)=\alpha\}$ is a unimodular level set of $\phi$; see [22, Proposition 11.2]. The expression in (3.2) is a finite sum since $\phi$ is a finite cover of $\mathbb{D}$, and in particular $\mathcal{C}_{\alpha}=\deg\phi$. The elements of $\mathcal{C}_{\alpha}$ are the eigenvalues of $U_{\alpha}$ and, in particular, the spectrum of the unitary $U_{\alpha}$ coincides precisely with the support of $\sigma_{\alpha}$, so that $\sigma(U_{\alpha})=\{\zeta\in\mathbb{T}\colon\phi(\zeta)=\alpha\}$.
@@ 3.2. Clark Measures for Two-Variable Inner Functions
A function $\phi\in H^{\infty}(\mathbb{D}^{2})$ is said to be inner if its non-tangential limit $\phi^{*}(\zeta)=\angle\lim_{z\to\zeta}\phi(z)$, which exists a.e. [34, Theorem 3.3.5], satisfies that $|\phi^{*}(\zeta)|=1$ at almost every $\mathbb{T}^{2}$. We say that $\phi$ is a rational inner function if $\phi=q/p$ for $q,p\in\mathbb{C}[z_{1},z_{2}]$, where $Z_{p}\cap\mathbb{D}^{2}=\emptyset$, and $q$ and $p$ are relatively prime. Polynomials with no zeros in $\mathbb{D}^{2}$ are called stable; note that $Z_{p}\cap\mathbb{T}^{2}\neq\emptyset$ is not ruled out, and in fact the case of boundary zeros is of considerable interest. One can show (see [34, Theorem 5.2.5]) that $\phi$ is a rational inner function in $\mathbb{D}^{2}$ if and only if
@@
Let us assume that $p$ is atoral, meaning in our context that $p$ and its reflection polynomial do not share a common factor (see [1] for a general discussion). In contrast to the one-variable setting, a rational inner function (RIF) in two variables can have singularities on $\mathbb{T}^{2}$: these occur at common zeros in $Z_{p}\cap Z_{\tilde{p}}\cap\mathbb{T}^{2}$. By a theorem of Knese ([27, Theorem C]; see also [10, Theorem 3.2] for an extension to general bounded rational functions), any RIF has non-tangential limits at every point of $\zeta\in\mathbb{T}^{2}$, and these non-tangential limits satisfy $\phi^{*}(\zeta)\in\mathbb{T}$. It should be noted that the function $\zeta\mapsto\phi^{*}(\zeta)$ is discontinuous on $\mathbb{T}^{2}$ whenever $Z_{p}\cap\mathbb{T}^{2}\neq\emptyset$.
@@
$(S^{j}_{\phi}f)(z)=(P_{K_{\phi}}T_{j}f)(z),\quad z\in\mathbb{D}^{2},$ (3.3)
where $P_{K_{\phi}}\colon H^{2}(\mathbb{D}^{2})\to K_{\phi}$ denotes the orthogonal projection of $H^{2}(\mathbb{D}^{2})$ onto its closed subspace $(\phi H^{2}(\mathbb{D}^{2}))^{\perp}$.
We turn to a discussion of Clark measures associated with a rational inner function $\phi$. For a fixed $\alpha\in\mathbb{T}$, we define the unimodular level set
$\mathcal{C}_{\alpha}=\overline{\{\zeta\in\mathbb{T}^{2}\colon\phi^{\ast}(\zeta)=\alpha\}}.$ (3.4)
As is explained in [11], we have the identification $\mathcal{C}_{\alpha}=\{\zeta\in\mathbb{T}^{2}\colon\tilde{p}(\zeta)-\alpha p(\zeta)=0\}$. A fortiori, the level set $\mathcal{C}_{\alpha}$ can be parameterized by analytic functions [11, Theorem 2.8] (see also the discussion in [4, Section 2], and in particular Lemma 2.5 therein).
@@
for some positive singular Borel measure on $\mathbb{T}^{2}$ and for one-variable Poisson kernels $P_{z_{j}}$. These measures $\{\sigma_{\alpha}\}$ are the Clark measures associated with $\phi$. Some basic properties of Clark measures in polydisks were established by Doubtsov in [20]; for instance, we have $\sigma_{\alpha_{1}}\perp\sigma_{\alpha_{2}}$ whenever $\alpha_{1}\neq\alpha_{2}$, as in one variable, and an analog of the Aleksandrov disintegration theorem holds for $\{\sigma_{\alpha}\}$ (see [20, Proposition 2.4]).
These measures admit a particularly concrete representation for $\phi$ rational inner. For each fixed $\zeta_{1}\in\mathbb{T}$, the function $z_{2}\mapsto\phi(\zeta_{1},z_{2})$ is a finite Blaschke product, with degree at most $n_{2}=\deg_{z_{2}}p$. The formula (3.2) describes the Clark measures associated with a finite Blaschke product, and from this one might hope to obtain a corresponding formula for two-variable RIFs by slicing. This is indeed possible, and one of the main results of [4]. Before we formulate this as a theorem, we need to introduce the notion of exceptional $\alpha\in\mathbb{T}$. We say that $\alpha\in\mathbb{T}$ is generic if $\mathcal{C}_{\alpha}$ does not contain any sets of the form $\{(\zeta,\eta)\colon\zeta\in\mathbb{T}\}$ or $\{(\eta,\zeta)\colon\zeta\in\mathbb{T}\}$ for some $\eta\in\mathbb{T}$. In other words, $\alpha$ is generic for $\phi$ if the corresponding level set does not contain any horizontal or vertical lines. If $\alpha$ fails to be generic, then we say that $\alpha$ is exceptional. Note that a RIF can have at most a finite number of exceptional $\alpha$ (see [4]).
@@
Let $\phi=\frac{\tilde{p}}{p}$ be a bidegree $(n_{1},n_{2})$ RIF, and let $\alpha\in\mathbb{T}$ be generic for $\phi$. Then, for $f\in C(\mathbb{T}^{2})$ the associated Clark measure $\sigma_{\alpha}$ satisfies
@@
For each $\alpha\in\mathbb{T}$, let the Clark embedding $J_{\alpha}:K_{\phi}\rightarrow L^{2}(\mu_{\alpha})$ be initially defined on the reproducing kernels $k^{\phi}_{w}$ by
$(J_{\alpha}k^{\phi}_{w})(z)=\frac{1-\alpha\overline{\phi(w)}}{(1-z_{1}\bar{w}_{1})(1-z_{2}\bar{w}_{2})}=(1-\alpha\overline{\phi(w)})k_{w}(z),$ (3.5)
for all $w\in\mathbb{D}^{2}$. Then the work of Doubtsov in [20, Theorem 3.1] shows that $J_{\alpha}$ extends to a linear isometry on $K_{\phi}$. In two or more variables, this isometry need not be surjective (viz. [20] and [9]), but fortunately, we have the following.
@@
Let $\phi=\frac{\tilde{p}}{p}$ be a bidegree $(n_{1},n_{2})$ RIF and let $\alpha\in\mathbb{T}$ be a generic value for $\phi$. Then the Clark embedding $J_{\alpha}\colon K_{\phi}\to L^{2}(\sigma_{\alpha})$ is unitary.
Corollary [4, Corollary 5.7] shows that for all but possibly finitely many $\alpha$, the densities appearing in the right-hand side of Theorem 3.3 are bounded functions on $\mathbb{T}$ which are non-zero Lebesgue almost everywhere. This implies that $L^{2}(\sigma_{\alpha})$ is infinite-dimensional. Together with Theorem 3.4, this implies the following fact, which is in contrast to the case of a one-variable finite Blaschke product.
@@
It is worth noting that this fact can also be deduced from other known results. For example, the existence of Agler decompositions for Schur functions on the bidisk immediately implies that for any nonconstant inner function $\phi$ on $\mathbb{D}^{2}$, the associated model space $K_{\phi}$ is infinite-dimensional.

### v2

@@ 2. A Simple Example
The family $\{\sigma_{\alpha}\}_{\alpha\in\mathbb{T}}$ of Clark measures associated with $\phi$ (formally defined in Subsection 3.2) can readily be identified, see e.g. [26, Example 4.2]. Namely, we have the pairing
$\int_{\mathbb{T}^{2}}f(\zeta)d\sigma_{\alpha}(\zeta)=\int_{\mathbb{T}}f(\zeta_{1},\alpha\bar{\zeta}_{1})|d\zeta_{1}|,\quad f\in C(\mathbb{T}^{2}).$
This shows $\mathrm{supp}\,\sigma_{\alpha}=\{(\zeta_{1},\alpha\bar{\zeta}_{1})\colon\zeta_{1}\in\mathbb{T}\}$. A collection of these support sets is graphed in Figure 1.
As is explained in [26], there is a surjective isometry $J_{\alpha}\colon K_{\phi}\to L^{2}(\sigma_{\alpha})$ for each value of $\alpha$. By [14, Theorem 1.3], this isometry acts by evaluation of $f\in K_{\phi}$ on $\mathrm{supp}\,\sigma_{\alpha}$, which is well-defined for $\sigma_{\alpha}$-a.e. $\zeta\in\mathbb{T}^{2}$. (In this example, one could also check directly that this evaluation map gives a unitary operator from $K_{\phi}\to L^{2}(\sigma_{\alpha})$.) Then, for almost every $\zeta\in\mathrm{supp}\,\sigma_{\alpha}$ and each power of $z_{j},$
$(J_{\alpha}U^{1}_{\alpha}z_{1}^{k})(\zeta)$ $=\zeta_{1}^{k+1}=\zeta_{1}\cdot\zeta_{1}^{k}=(M_{\zeta_{1}}J_{\alpha}z_{1}^{k})(\zeta)$
$(J_{\alpha}U^{1}_{\alpha}z_{2}^{l})(\zeta)$ $=\alpha(\alpha\bar{\zeta_{1}})^{l-1}=\zeta_{1}\cdot(\alpha\bar{\zeta}_{1})^{l}=(M_{\zeta_{1}}J_{\alpha}z_{2}^{l})(\zeta),$
so that $J_{\alpha}U^{1}_{\alpha}=M_{\zeta_{1}}J_{\alpha}$, where $M_{\zeta_{1}}\colon L^{2}(\sigma_{\alpha})\to L^{2}(\sigma_{\alpha})$ is the unitary furnished by multiplication by the coordinate function $\zeta_{1}$. Similarly,
$(J_{\alpha}U^{2}_{\alpha})(z_{1}^{k})=\alpha\bar{\zeta}_{1}\cdot\zeta_{1}^{k}\quad\textrm{and}\quad(J_{\alpha}U^{2}_{\alpha})(z_{2}^{l})=\alpha\bar{\zeta}_{1}\cdot(\alpha\bar{\zeta}_{1})^{l},$
which implies that $J_{\alpha}U^{2}_{\alpha}=M_{\alpha\bar{\zeta}_{1}}J_{\alpha}=M_{\zeta_{2}}J_{\alpha}$.
Finally, one can check directly (or using Theorem 5.1 below) that the Taylor joint spectrum of $(U^{1}_{\alpha},U^{2}_{\alpha})$ (see Subsection 3.4 for a definition) is
@@ 3.1. Clark Theory in One Variable
Recall that each inner function $\phi\colon\mathbb{D}\to\mathbb{C}$ gives rise to a family of Clark measures $\{\sigma_{\alpha}\}_{\alpha\in\mathbb{T}}$ via (1.1); each $\sigma_{\alpha}$ is singular with respect to Lebesgue measure on the unit circle, and moreover $\sigma_{\alpha_{1}}\perp\sigma_{\alpha_{2}}$ for distinct $\alpha_{j}$. Clark measures can be identified as spectral measures associated with certain unitary operators acting on the model space $K_{\phi}=H^{2}(\mathbb{D})\ominus\phi H^{2}(\mathbb{D})$. Namely, while the compressed shift $S_{\phi}\colon K_{\phi}\to K_{\phi}$ is not generally unitary (see eg. [28, Chapter 9]), they do admit one-dimensional perturbations to unitaries. More specifically, for each $\alpha\in\mathbb{T}$, one can define the unitary operator $U_{\alpha}$ by
(3.1) $U_{\alpha}f:=S_{\phi}f+\frac{\alpha}{1-\overline{\phi(0)}\alpha}\langle f,Ck^{\phi}_{0}\rangle k^{\phi}_{0},\quad f\in K_{\phi},$
where $k^{\phi}_{0}$ is the reproducing kernel for $K_{\phi}$ at the origin, and $Ck^{\phi}_{0}$ is its conjugation. If $\phi(0)=0$, then $k^{\phi}_{0}=1$ and (3.1) simplifies to
@@
and then extending to $K_{\phi}$ by linearity. Then the following theorem is well known.
@@
See [28, Chapter 8, Chapter 11] for a detailed exposition, including proofs of the above statements.
@@
Then the associated model space $K_{\phi}$ has $\dim K_{\phi}=\deg\phi$ (see [28, Proposition 5.16]), hence is finite-dimensional, and is spanned by reproducing kernels and their derivatives. In fact, $K_{\phi}$ is finite dimensional precisely when $\phi$ is a finite Blaschke product.
@@
(3.2) $\sigma_{\alpha}=\sum_{\zeta\in\mathcal{C}_{\alpha}}\frac{1}{|\phi^{\prime}(\zeta)|}\delta_{\zeta},$
where $\delta_{\zeta}$ denotes point evaluation at $\zeta$ and $\mathcal{C}_{\alpha}=\{\zeta\in\mathbb{T}\colon\phi(\zeta)=\alpha\}$ is a unimodular level set of $\phi$; see [28, Proposition 11.2]. The expression in (3.2) is a finite sum since $\phi$ is a finite cover of $\mathbb{D}$, and in particular $\mathcal{C}_{\alpha}=\deg\phi$. The elements of $\mathcal{C}_{\alpha}$ are the eigenvalues of $U_{\alpha}$ and, in particular, the spectrum of the unitary $U_{\alpha}$ coincides precisely with the support of $\sigma_{\alpha}$, so that $\sigma(U_{\alpha})=\{\zeta\in\mathbb{T}\colon\phi(\zeta)=\alpha\}$.
@@ 3.2. Clark Measures for Two-Variable Inner Functions
A function $\phi\in H^{\infty}(\mathbb{D}^{2})$ is said to be inner if its non-tangential limit $\phi^{*}(\zeta)=\angle\lim_{z\to\zeta}\phi(z)$, which exists a.e. [41, Theorem 3.3.5], satisfies that $|\phi^{*}(\zeta)|=1$ at almost every $\mathbb{T}^{2}$. We say that $\phi$ is a rational inner function if $\phi=q/p$ for $q,p\in\mathbb{C}[z_{1},z_{2}]$, where $Z_{p}\cap\mathbb{D}^{2}=\emptyset$, and $q$ and $p$ are relatively prime. Polynomials with no zeros in $\mathbb{D}^{2}$ are called stable; note that $Z_{p}\cap\mathbb{T}^{2}\neq\emptyset$ is not ruled out, and in fact the case of boundary zeros is of considerable interest. One can show (see [41, Theorem 5.2.5]) that $\phi$ is a rational inner function in $\mathbb{D}^{2}$ if and only if
@@
Let us assume that $p$ is atoral, meaning in our context that $p$ and its reflection polynomial do not share a common factor (see [2] for a general discussion). If $p$ is stable and atoral, then $p$ also has no zeros on the faces of the bidisk $\mathbb{D}\times\mathbb{T}$ and $\mathbb{T}\times\mathbb{D}$, see [33, Lemma 10.1]. In contrast to the one-variable setting, a rational inner function (RIF) in two variables can have singularities on $\mathbb{T}^{2}$: these occur at common zeros in $Z_{p}\cap Z_{\tilde{p}}\cap\mathbb{T}^{2}$. By a theorem of Knese ([33, Theorem C]; see also [15, Theorem 3.2] for an extension to general bounded rational functions), any RIF has non-tangential limits at every point of $\zeta\in\mathbb{T}^{2}$, and these non-tangential limits satisfy $\phi^{*}(\zeta)\in\mathbb{T}$. It should be noted that the function $\zeta\mapsto\phi^{*}(\zeta)$ is discontinuous on $\mathbb{T}^{2}$ whenever $Z_{p}\cap\mathbb{T}^{2}\neq\emptyset$ (see for instance [39]).
@@
(3.3) $(S^{j}_{\phi}f)(z)=(P_{{\phi}}T_{j}f)(z),\quad z\in\mathbb{D}^{2},$
where $P_{{\phi}}\colon H^{2}(\mathbb{D}^{2})\to K_{\phi}$ denotes the orthogonal projection of $H^{2}(\mathbb{D}^{2})$ onto its closed subspace $(\phi H^{2}(\mathbb{D}^{2}))^{\perp}$. As is proved in [9, Theorem 1.1], a closed subspace $\mathcal{Q}\subset H^{2}(\mathbb{D}^{2})$ invariant under $(T_{1}^{*},T_{2}^{*})$ is in fact of the form $\mathcal{Q}=K_{\phi}$ for some inner function $\phi$ precisely if
$(I-(P_{\mathcal{Q}}T_{j})^{*}P_{\mathcal{Q}}T_{j})(I-P_{\mathcal{Q}}T_{k}(P_{\mathcal{Q}}T_{k})^{*})=0,\quad j\neq k,$
where $I$ is the identity operator and $P_{\mathcal{Q}}\colon H^{2}(\mathbb{D}^{2})\to\mathcal{Q}$ denotes orthogonal projection; this result remains valid in $d$ variables. Bhattacharjee, Das, Debnath, and Sarkar further provide a characterization of $d$-tuples of commuting operators which are unitarily equivalent to compressions of the shift to a model space [9, Theorem 3.2]. Thus, many of our results apply to pairs of operators covered by this theorem.
We turn to a discussion of Clark measures associated with a rational inner function $\phi$. For a fixed $\alpha\in\mathbb{T}$, define the unimodular level set
(3.4) $\mathcal{C}_{\alpha}=\overline{\{\zeta\in\mathbb{T}^{2}\colon\phi^{\ast}(\zeta)=\alpha\}}.$
As is explained in [17], when $\phi=\tfrac{\tilde{p}}{p}$, we have the identification $\mathcal{C}_{\alpha}=\{\zeta\in\mathbb{T}^{2}\colon\tilde{p}(\zeta)-\alpha p(\zeta)=0\}$. A fortiori, the level set $\mathcal{C}_{\alpha}$ can be parameterized by analytic functions [17, Theorem 2.8] (see also the discussion in [6, Section 2], and in particular Lemma 2.5 therein).
@@
for some positive singular Borel measure $\sigma_{\alpha}$ on $\mathbb{T}^{2}$ and for one-variable Poisson kernels $P_{z_{j}}$. These measures $\{\sigma_{\alpha}\}$ are the Clark measures associated with $\phi$. Some basic properties of Clark measures in polydisks were established by Doubtsov in [26]; for instance, we have $\sigma_{\alpha_{1}}\perp\sigma_{\alpha_{2}}$ whenever $\alpha_{1}\neq\alpha_{2}$, as in one variable, and an analog of the Aleksandrov disintegration theorem holds for $\{\sigma_{\alpha}\}$ (see [26, Proposition 2.4]).
These measures admit a particularly concrete representation for $\phi$ rational inner. For each fixed $\zeta_{1}\in\mathbb{T}$, the function $z_{2}\mapsto\phi(\zeta_{1},z_{2})$ is a finite Blaschke product, with degree at most $n_{2}=\deg_{z_{2}}p$. The formula (3.2) describes the Clark measures associated with a finite Blascke product, and from this one might hope to obtain a corresponding formula for two-variable RIFs by slicing. This is indeed possible, and is one of the main results of [6]. Before we formulate this as a theorem, we need to introduce the notion of exceptional $\alpha\in\mathbb{T}$. We say that $\alpha\in\mathbb{T}$ is generic if $\mathcal{C}_{\alpha}$ does not contain any sets of the form $\{(\zeta,\eta)\colon\zeta\in\mathbb{T}\}$ or $\{(\eta,\zeta)\colon\zeta\in\mathbb{T}\}$ for some $\eta\in\mathbb{T}$. In other words, $\alpha$ is generic for $\phi$ if the corresponding level set does not contain any horizontal or vertical lines. If $\alpha$ fails to be generic, then we say that $\alpha$ is exceptional. Note that a RIF of two variables can have at most a finite number of exceptional $\alpha$ (see [6]).
@@
Let $\phi=\frac{\tilde{p}}{p}$ be a bidegree $(n_{1},n_{2})$ RIF with $n_{1},n_{2}\geq 1$, and let $\alpha\in\mathbb{T}$ be generic for $\phi$. Then for $f\in C(\mathbb{T}^{2})$, the associated Clark measure $\sigma_{\alpha}$ satisfies
@@
For each $\alpha\in\mathbb{T}$, let the Clark embedding $J_{\alpha}:K_{\phi}\rightarrow L^{2}(\sigma_{\alpha})$ be initially defined on the reproducing kernels $k^{\phi}_{w}$ by
(3.5) $(J_{\alpha}k^{\phi}_{w})(z)=\frac{1-\alpha\overline{\phi(w)}}{(1-z_{1}\bar{w}_{1})(1-z_{2}\bar{w}_{2})}=(1-\alpha\overline{\phi(w)})k_{w}(z),$
for all $w\in\mathbb{D}^{2}$. Then the work of Doubtsov in [26, Theorem 3.1] shows that $J_{\alpha}$ extends to a linear isometry on $K_{\phi}$. In two or more variables, this isometry need not be surjective (viz. [26] and [14]), but fortunately, we have the following.
@@
Let $\phi$ be a bidegree $(n_{1},n_{2})$ RIF with $n_{1},n_{2}\geq 1$ and let $\alpha\in\mathbb{T}$ be a generic value for $\phi$. Then the Clark embedding $J_{\alpha}\colon K_{\phi}\to L^{2}(\sigma_{\alpha})$ is unitary.
As a note, even though Theorem 4.2 is proved in [6] for RIFs of the form $\frac{\tilde{p}}{p}$, the proof extends without issue to RIFs with an additional monomial factor. Corollary [6, Corollary 5.7] shows that for all but possibly finitely many $\alpha$, the densities appearing in the right-hand side of Theorem 3.3 are bounded functions on $\mathbb{T}$ which are non-zero Lebesgue almost everywhere. This implies that $L^{2}(\sigma_{\alpha})$ is infinite-dimensional. Together with Theorem 3.4, this implies the following fact, which is in contrast to the case of a one-variable finite Blaschke product.
@@
3.3. Agler decompositions
The conclusion in Corollary 3.5 can also be deduced from other known results. For example, the existence of Agler decompositions for Schur functions on the bidisk (see [1, 5]) immediately implies that for any nonconstant inner function $\phi$ on $\mathbb{D}^{2}$, the associated model space $K_{\phi}$ is infinite-dimensional.
As is explained in [13, 12], the model space $K_{\phi}$ can be decomposed as
(3.6) $K_{\phi}=S_{1}^{\mathrm{max}}\oplus S^{\mathrm{min}}_{2},$
where the subspace $S_{1}^{\mathrm{max}}$ is the largest subspace of $K_{\phi}$ invariant with respect to $T_{1}$. It is also true (but non-trivial to prove) that $S_{2}^{\mathrm{min}}$ is $T_{2}$-invariant. Moreover, we have
$S_{1}^{\mathrm{max}}=\mathcal{H}\left(\frac{K_{1}(z,w)}{1-\bar{w}_{1}z_{1}}\right)\ \ \text{ and }\ \ S_{2}^{\mathrm{min}}=\mathcal{H}\left(\frac{K_{2}(z,w)}{1-\bar{w}_{2}z_{2}}\right),$
for positive kernels $K_{1},K_{2}:\mathbb{D}^{2}\times\mathbb{D}^{2}\rightarrow\mathbb{C}$, where
$\mathcal{H}(K_{1})=S_{1}^{\mathrm{max}}\ominus T_{1}S_{1}^{\mathrm{max}}\ \ \text{ and }\ \ \mathcal{H}(K_{2})=S_{2}^{\mathrm{min}}\ominus T_{2}S_{2}^{\mathrm{min}}$
and thus,
$S_{1}^{\mathrm{max}}=\bigoplus_{k=0}^{\infty}z_{1}^{k}\mathcal{H}(K_{1})\ \ \text{ and }\ \ S_{2}^{\mathrm{min}}=\bigoplus_{l=0}^{\infty}z_{2}^{l}\mathcal{H}(K_{2}).$
We also recall (see [13, Theorem 3.2]) that if a RIF $\phi$ satisfies $\deg_{z_{1}}\phi=n_{1}$ and $\deg_{z_{2}}\phi=n_{2}$, then
(3.7) $\dim(\mathcal{H}(K_{1}))=n_{2}\quad\textrm{and}\quad\dim(\mathcal{H}(K_{2}))=n_{1}.$